OpenStudy (anonymous):

A certain job can be done by 72 men in 100 days. There were 80 men at the start of the project but after 40 days, 30 of them had to be transferred to another project. How long will it take the remaining workforce to complete the job?

OpenStudy (anonymous):

i get 80 days but i could be wrong

OpenStudy (anonymous):

yes the answer is 80,but how come, i really forgot how to solve work problem

OpenStudy (anonymous):

there is probably a snappy way to do this, but i don't know of one. i said "if 72 men take 100 days to complete 1 job, then each man completes \[\frac{1}{7200}\]of the job per day"

OpenStudy (anonymous):

you had 80 men working for 40 days and \[\frac{80\times 40}{7200}=\frac{4}{9}\] meaning they completed 4/9 th of the job, leaving 5/9 to be completed.

OpenStudy (anonymous):

now 30 leave, so there are 50 men left, and we need to solve \[\frac{50\times x}{7200}=\frac{5}{9}\] and that gives \[x=80\]

OpenStudy (anonymous):

like i said, there is probably a snappier way to do it

OpenStudy (anonymous):

i should recall to solve this kind of problem, thanks then :D